OpenStudy (raffle_snaffle):

Cushing, Inc. invested $226821 in an account paying 7% per year, compounded monthly. How much will Cushing have in its account in 9 years?

OpenStudy (raffle_snaffle):

I have a solution. I would like for someone to look at it plz.

OpenStudy (raffle_snaffle):

@hartnn

OpenStudy (raffle_snaffle):

This is how I solved it.

OpenStudy (raffle_snaffle):

r = 7/12 = 58.33% per month this is the nominal annual interest rate

OpenStudy (raffle_snaffle):

then I used i =(1+(0.5833/12))^12-1 = 76.75% this is the effective annual interest rate

OpenStudy (raffle_snaffle):

Then I used F =P(1+i)^n = 226821(1+0.7675)^9 = 38198334.03

OpenStudy (raffle_snaffle):

is it right?

OpenStudy (raffle_snaffle):

i did it wrong... I think i see something wrong

OpenStudy (raffle_snaffle):

Can I use F = P(e^rn)?

OpenStudy (raffle_snaffle):

cant use that...

OpenStudy (raffle_snaffle):

I need to find i = r/m right then I can use F = P(F/P, i, n)??

OpenStudy (raffle_snaffle):

I got F = 239,010.77 as my answer

OpenStudy (raffle_snaffle):

hello????

hartnn (hartnn):

for monthly compounded problems, I generally use this formula, \(F= P (1+i/12)^{n/12}\) P = 226821 i = 0.07 n= 9

OpenStudy (raffle_snaffle):

what?? why is n/12?

hartnn (hartnn):

oops, thats 12n and not n/12

hartnn (hartnn):

\(\large F= P (1+i/12)^{12n}\)

hartnn (hartnn):

and i got 425102.6949...

OpenStudy (raffle_snaffle):

why is it 12n?

OpenStudy (raffle_snaffle):

let me do something really quick.

hartnn (hartnn):

for 1st month, (1+i/12)^n for 2nd month (1+i/12)^n and so on, for 12th month too (1+i/12)^n multiplying all these together, \( [(1+i/12)^n]^{12}\)

OpenStudy (raffle_snaffle):

I am not understanding what r is in my text r = i*m which is your nominal annual interest rate and i = r/m is your effective annual interest rate

OpenStudy (raffle_snaffle):

so it's already a nominal interest rate?

hartnn (hartnn):

i am not familiar with those terms i can explain you how we have i/12 and 12n interest rate per year = 7% so interest rate per month = 7%/12 and interest is compounded monthly, so there will be 12 such compounds (1+i/12)^n, one for each month, multiplied together

hartnn (hartnn):

similarly, if we had to compound daily, the formula would become, \(\Large F = P (1+i/365)^{365n}\)

OpenStudy (raffle_snaffle):

I am going to try to solve this using another method and use your solution as the key. Thanks for you rhelp

hartnn (hartnn):

yeah, verify if you get 425102.6949... from other method.. welcome ^_^